OpenStudy (anonymous):

how do you simplify this ratio: 12cm/4m?

OpenStudy (bahrom7893):

3cm jutst divide.. hello??

OpenStudy (anonymous):

look at the units bahrom

OpenStudy (bahrom7893):

OHHH LOL sorry guys

OpenStudy (anonymous):

12cm /400 cm

OpenStudy (anonymous):

it's ok =P

OpenStudy (anonymous):

Both 12 and 4 are divisible by 4, so just divide them both by 4.\[\frac{12 \times \frac{1}{3}}{4 \times \frac{1}{3}} = \frac{3}{1} = 3\]

OpenStudy (anonymous):

the same for you alex. Look at the units first

OpenStudy (anonymous):

Oh oops, I didn't realize they were different units :P